OpenStudy (goformit100):

Determine all positive integers n such that 5^n−1 can be written as a product of an even number of consecutive integers.

OpenStudy (goformit100):

@tanjeetsarkar96

OpenStudy (anonymous):

wait m thinking ..!!

OpenStudy (anonymous):

is the total n-1 raised to power ??? @goformit100

OpenStudy (goformit100):

ok

OpenStudy (goformit100):

@Mandila Help me Sister

OpenStudy (anonymous):

@tanjeetsarkar96 I think it would be\[5^{n}-1\]since that is always even and the product of an even number of integers is also always even, while \[5^{n-1}\]is only even if n=1.

OpenStudy (goformit100):

Ok...

OpenStudy (anonymous):

as n must be even let n belong to integers then 2n will belong to even nos as soln will be consecutive nos equation becomes 2*n(2*(n+1))=5^n-1 2n(2n+2)=5^n-1 4n^2+4n=5^n-1 n^2+n-(5^n-1)/4=0 solving quadratic we get n=-1+-(\[\sqrt{1^2-(-[5^n-1]/4)}\] put n=1 you get 2=[-1+-{5^0.5}]/2 which is not soln put n=2 you get 4=[-1+-{5^1}]/2 which gives one extraneous soln and other correct soln put n=3 you get 6=[-1+-{5^1.5}]/2 which is not soln put values gap gets wider and wider therefore solution is 2 only.....

OpenStudy (anonymous):

@go formit is my answer ryt

OpenStudy (anonymous):

im like clueless here! my mind is filled with too much econ i guess!

OpenStudy (anonymous):

@mandila pls go thru my answer is it ryt

OpenStudy (anonymous):

(Y)

OpenStudy (anonymous):

ty @Mandila

OpenStudy (anonymous):

question gives a different meanin! it has to be product of 2 consecutive even numbers! i was soo confused! otherwise solving this would be like magic! :)

OpenStudy (loser66):

to me, it cannot be done. product of two even consecutive number is an even number let 2x is the first one, (it is an even, so, it must be formed as 2x ) and the second one is 2x+2 the product is 2x*(2x+2) = 4 (x)(x+1) is an even number whereas 5 ^(n-1) is odd numbers with all values of n how can an even number = an odd number?

OpenStudy (anonymous):

-1 is not in the index part right?

OpenStudy (loser66):

@Mandila you ask me or the asker?

OpenStudy (loser66):

if n is odd--> n-1 is even if n is even--> n -1 is odd. however, no matter what it is, 5^ ( odd) is an odd, and 5^(even) is an odd, too

OpenStudy (anonymous):

asker

OpenStudy (loser66):

@goformit100 you can put it in logic by \[5^1 = 5 mod 10\] \[5^2 =5 mod 10\] \[5^3 =5 mod 10\] ................................ \[5^{n-1}=5 mod 10\] so, the last digit of 5^(n-1) is 5 , that shows it is an odd number that logic for the first part, the second part is even part as I show above. then combine them to give out conclusion.

OpenStudy (goformit100):

Thankyou

OpenStudy (loser66):

yw